OpenStudy (anonymous):

What are the factors of x2 - 81? (x + 9)(x - 9) (x - 9)(x - 9) (x + 3)(x - 27) Prime

Nnesha (nnesha):

difference of square s \[\huge\rm a^2 - b^2 = (a+b)(a-b)\]

pooja195 (pooja195):

Diffrence of squares \[\huge~\rm~\sqrt{ x^2}=\sqrt{81}\]

Nnesha (nnesha):

take square root of both term write your answer in two parenthses (sqrt of first term + sqrt of 2nd term )( sqrt of first term - sqt of 2nd term )

OpenStudy (anonymous):

You can use foil here & just work backwards aswell, so looking at answer choice #1, (x+9)(x-9) "F" first: X * X = X^2 "O" outside: X * 9 = 9x "I" inside: -9 * X = -9x "L" last: 9 * (-9) = -81 So you have: x^2 + 9x - 9x -81 Now simplify: x^2 -81 Thats how i solved my equations in school. hope it helps :)

OpenStudy (anonymous):

help plz @Mehek14

Mehek (mehek14):

what do you think is the answer?

OpenStudy (anonymous):

B?

Mehek (mehek14):

no -9 * -9 = +81 not -81

Mehek (mehek14):

it would be (x + 9)(x - 9)

Mehek (mehek14):

because x*x = \(x^2\) \(9x+-9x=0\\9*-9=-81\\x^2+0-81\\x^2-81\)

OpenStudy (anonymous):

oh ok i think i get it now